# Conformal mapping Coordinates Physics-Informed Neural Networks (CoCo-PINNs): learning neural networks for designing neutral inclusions

### v1

@@ D.2 Experimental setup and parameters
In common settings for experiments, we set $\sigma_{m}=1,\sigma_{m}=5$ and $\sigma_{c}=3,4,6,7$ are also used. The conformal radius $\gamma=1$ and $L=5$. which determines the domain of the conformal map, and sample points are located in the square box of $[-5,5]^{2}$. We generated collocation points $|\Omega^{\text{ext}}|=21,808,|\Omega^{\text{int}}|=6,000$, and $|\partial\Omega|=6,000$ with a single epoch so that collocation points are all fixed. For the neural networks’ architecture, we use 4 hidden layers with a width of 20 and the Tanh activation due to its smoothness. We use Adam optimizers under 25,000 iterations adjusting learning rates lr pinn: all types of neural networks and lr inv: for interface (inverse) parameters decaying with $\eta$ per 1000 iterations along the learning rate schedulers. We fix both learning rates as $10^{-3}$ and $\eta=0.7$. We use the CoCo-PINNs with $p=p^{(20)}$ which takes a real number $p_{0}$ and complex numbers $\{p_{i}\}_{i=1}^{20}$ as inverse parameters. To verify the classical PINNs’ result, we use the Fourier fitting with an order of $20$. The Fourier fitting is explained in Section D.4. Note that using too high order makes singularities, and hence, it represents awful credibility. We set the initial value of the interface function to 5 for the square and spike shape to make it easier to satisfy the condition of being a positive function.
@@
Remark on the environments  Traditionally, hyperparameter tuning is important to acquire the PINNs’ performance. Since balancing them is much more complicated due to the difficulty of the PDEs, we need to choose the appropriate values to enhance PINNs. Here are the brief guidelines for our settings: 1) When we balance the number of collocation points, although the exterior domain is much larger than the interior, fitting the background fields on the exterior domain has been designed to be less complex. So we set 4 times of interior points. Also, the boundary effect is essential for neutral inclusion, we put more points on the boundary same with the interior’s points. 2) Re-sampling collocation points can lead to the uncertainty of the interface function. If we sample again for each epoch, PINNs are hardly adapted for new sample points, especially for the boundary condition loss, which is quite important for the neutral inclusion effects. It is also from the experimental results that the results of re-sampling are worse than those of fixed-sampling. Hence, we fix the points and proceed to a single epoch. 3) To control the learning rates adaptively, we use the Adam optimizers so that we can handle the sensitivity of Fourier coefficients much better. Since Adam has faster convergence and robustness, they can adjust the learning rates appropriately during the training.
@@
11:S = Numeric[$\left[\{p_{k}\}_{k=1}^{n}\right]$
12:$u_{p}=\mbox{Re}\left[H+Sw\right]\cdots\cdots$eq. 2.3.
@@
D.3 Conformal maps for various shapes
The shapes shown in Figure 4.1 are defined by the conformal map $\Psi(w):\{w\in\mathbb{C}:|w|>1\}\to\mathbb{C}\setminus\overline{\Omega}$ as follows:
$\Psi(w)$ $=w+\frac{1}{10w^{3}},$ (D.1)
$\Psi(w)$ $=w+\frac{1}{4w}+\frac{1}{8w^{2}}+\frac{1}{10w^{3}},$ (D.2)
$\Psi(w)$ $=w+\frac{1}{10w}+\frac{1}{4w^{2}}-\frac{1}{20w^{3}}+\frac{1}{20w^{4}}-\frac{1}{25w^{5}}+\frac{1}{50w^{6}},$ (D.3)
$\Psi(w)$ $=w-\frac{1}{10w^{9}}.$ (D.4)
The eqs. D.1, D.2, D.3 and D.4 present ‘square’, ‘fish’, ‘kite’, and ‘spike’, respectively.
D.4 Fourier fittings
In order to ascertain whether the trained forward solution is true or not, it is necessary to identify the Fourier series that is sufficiently similar to the original interface function and, hence, achieve the real analytic solution. We utilize the Fourier series approximation for each interface function. Figure D.1 presents the difference between the interface function and the Fourier series we used. We denote $p_{\text{F}}$ as the Fourier series corresponding to the $p_{\text{NN}}$. Given that $p_{\text{NN}}$ is sufficiently close to $p_{\text{F}}$, it is reasonable to utilize the analytic solution obtained by $p_{\text{F}}$ in order to ascertain the credibility of the classical PINNs results.
@@ D.4 Fourier fittings
The relative errors and the Fourier fittings for each shape are given by Tables 5 and D.1, respectively.
@@
Figure D.1: Fourier fitting the interface function for each domain
D.5 Neutral inclusion effect for arbitrary fields
@@
In this subsection, we present the neutral inclusion effect of shapes and different fields $H(x_{1},x_{2})=x_{1},x_{2}$, and $2x_{1}-x_{2}$. The neutral inclusion effects for one random experiment results are given in Table 6.
@@
CoCo- PINNs square 5.004e-03 8.164e-03 2.801e-02
fish 4.050e-04 2.694e-04 1.868e-03
kite 4.358e-04 1.654e-04 1.928e-03
spike 2.230e-03 5.847e-03 1.498e-02
After many fair experiments with both CoCo-PINNs and classical PINNs, we concluded that CoCo-PINNs were superior. After that, we tested the neutral inclusion effect for each shape by utilizing the CoCo-PINNs. Figure D.2 presents the results. In the case of unit circle inclusion, exact neutral inclusion appeared. The shape we used may have no interface functions that make exact neutral inclusions. Notwithstanding, the CoCo-PINNs results for the neutral inclusion effect are, to some extent, satisfactory.
Figure D.2: Neutral inclusion effects for various shapes and background fields.
D.6 Illustrations for credibility
All experiments described in Table 1 are illustrated in Figures D.4 and D.3, by utilizing CoCo-PINNs and classical PINNs, respectively. We illustrate the pairs of
@@ D.6 Illustrations for credibility
for each shape in Figure 4.1.
Figure D.3: Experiment results by using the classical PINNs
Figure D.4: Experiment results by using the CoCo-PINNs
As shown Figures D.4 and D.3, the trained forward solutions by CoCo-PINNs $u_{\text{NN}}$ are more similar to analytic solution $u_{p}$ using the interface parameter given by the CoCo-PINNs’ training results than classical PINNs’ one.

### v2

@@ 4.1.2 Experimental setup and parameters
In the experimental settings, we set $\sigma_{m}=1,\sigma_{c}=5$ and additional values $\sigma_{c}=3,4,6,7$ are considered. The conformal radius $\gamma=1$ and $L=5$, which determines the domain of the conformal map. Sample points are located within the square $5,5]^{2}$. Collocation points $|\Omega^{\text{ext}}|=21,808,|\Omega^{\text{int}}|=6,000$, and $|\partial\Omega|=6,000$ were generated once(i.e. fixed throughout the training). We employ neural networks with four hidden layers of width 20 and use the Tanh activation function for its smoothness. Training proceeds for 25,000 iterations with the Adam optimizers under adjusting learning rates lr pinn: all types of neural networks, and lr inv: for interface (inverse) parameters decaying with $\eta$ per 1000 iterations along the learning rate schedulers. We fix both learning rates as $10^{-3}$ and $\eta=0.8$. In the CoCo-PINNs’ structure with $p=p^{(20)}$ which takes a real number $p_{0}$ and complex numbers $\{p_{i}\}_{i=1}^{20}$ as inverse parameters. To validate the classical PINNs’ result, we use the Fourier fitting with an order of $20$. The Fourier fitting is explained in Section 4.1.4. Note that using higher-order terms may introduce singularities, potentially degrading the credibility of the solution. For the square and spike-shaped inclusions, the interface function was initialized with a value of 5 to help satisfy the positivity condition more easily.
@@
Remark on the environments  Hyperparameter tuning is important for achieving the PINNs’ performance. Since balancing them is much more complicated due to the difficulty of the PDEs, we need to choose the appropriate values to enhance PINNs. Here are the brief guidelines for our settings: 1) Collocation Points: Although the exterior domain is significantly larger than the interior, fitting the background fields in the exterior region is designed to be less complex. The number of collocation points in the exterior domain is set to be four times that of the interior. Since the boundary behavior is crucial for achieving neutral inclusion, the number of boundary points is set equal to that of the interior. 2) Fixed Sampling vs. Re-sampling Re-sampling collocation points can lead to the uncertainty of the interface function. When points are re-sampled, PINNs often struggle to adapt to the new locations, particularly in enforcing boundary condition losses, which are critical for achieving accurate neutral inclusion effects. Empirical results demonstrate that fixed sampling yields more reliable outcomes; therefore, collocation points remain constant throughout training. 3) Adaptive Learning Rates To control the learning rates adaptively, we use the Adam optimizers so that we can handle the sensitivity of Fourier coefficients much better. Since Adam has faster convergence and robustness, thereby improving stability and accuracy.
@@
11:  Compute $u_{p}$ from $\{p_{k}\}_{k=1}^{n}$ using the analytic solution derived in [45].
@@
4.1.3 Conformal maps for various shapes
The shapes shown in Figure 4 are defined by the conformal map $\Psi(w):\{w\in\mathbb{C}:|w|>1\}\to\mathbb{C}\setminus\overline{\Omega}$ as follows:
$\Psi(w)$ $=w+\frac{1}{10w^{3}},$ (13)
$\Psi(w)$ $=w+\frac{1}{4w}+\frac{1}{8w^{2}}+\frac{1}{10w^{3}},$ (14)
$\Psi(w)$ $=w+\frac{1}{10w}+\frac{1}{4w^{2}}-\frac{1}{20w^{3}}+\frac{1}{20w^{4}}-\frac{1}{25w^{5}}+\frac{1}{50w^{6}},$ (15)
$\Psi(w)$ $=w-\frac{1}{10w^{9}}.$ (16)
The 13, 14, 15 and 16 present ‘square’, ‘fish’, ‘kite’, and ‘spike’, respectively.
4.1.4 Fourier fittings
In order to ascertain whether the trained forward solution is true or not, it is necessary to identify the Fourier series that is sufficiently similar to the original interface function and, hence, achieve the real analytic solution. We utilize the Fourier series approximation for each interface function. Figure 5 presents the difference between the interface function and the Fourier series we used. We denote $p_{\text{F}}$ as the Fourier series corresponding to the $p_{\text{NN}}$. Given that $p_{\text{NN}}$ is sufficiently close to $p_{\text{F}}$, it is reasonable to utilize the analytic solution obtained by $p_{\text{F}}$ in order to ascertain the credibility of the classical PINNs results.
@@ 4.1.4 Fourier fittings
The relative errors and the Fourier fittings for each shape are given by Tables 1 and 5, respectively.
@@
Figure 5: Fourier fitting: The interface function obtained by classical PINNs and its Fourier series approximation.
4.2 Neutral inclusion
We present experimental results demonstrating the successful achievement of the neutral inclusion effect using CoCo-PINNs. For training, we use two background solutions, $H_{u}(x)=x_{1}$ and $H_{v}(x)=x_{2}$. We then demonstrate the neutral inclusion effect using three background fields: $H(x)=x_{1}$, $x_{2}$, and $2x_{1}-x_{2}$; see Theorem 3.1 for the theoretical justification.
Inclusions generally yield perturbations in the applied background fields. However, the domain $\Omega$, with the interface function $p(x)$ trained from CoCo-PINNs, achieves the neutral inclusion effects across all three test background fields, as shown by the level curves of the analytic solutions $u_{p}$ in Figure 6.
Figure 6: Neutral inclusion effect after training: For the square’, fish’, kite’, spike’ shaped inclusions, the interface function $p^{(n)}$ is simultaneously trained using the two background solutions $H_{u}(x)=x_{1}$ and $H_{v}(x)=x_{2}$ through the CoCo-PINNs framework.
4.3 Credibility of classical PINNs and CoCo-PINNs
Table 2: Credibility results.
Shape Credibility
Classical PINNs square 2.843e-03 1.371e-01
fish 4.432e-03 1.546e-01
kite 1.246e-03 2.483e-01
spike 8.798e-03 4.301e-01
CoCo- PINNs square 2.802e-03 9.244e-02
fish 7.468e-04 1.067e-01
kite 4.299e-04 2.280e-01
spike 9.320e-03 4.038e-01
We investigate the credibility of the two methods. We examine whether the exterior part of the trained forward solution $u^{\text{ext}}_{\text{NN}}$ matches the analytic solution $u_{p}$. Recall that we denote $u_{p}$ as the analytic solution when the coefficients of the interface function $p$ are given by training. Once training is complete, CoCo-PINNs provide the expansion coefficients of the interface function directly. For classical PINNs, where the interface function is represented by neural networks, we compute the Fourier coefficients of $p_{\text{NN}}$. We use the Fourier series expansion up to a sufficiently high order to ensure that the difference between the neural network-designed interface function and its Fourier series is small (see Figure 5 in Section 4.1.4). We conduct experiments with four inclusion shapes, as shown in Figure 4, and the corresponding results are presented subsequently in Figures 10 and 11 (Section 5.2).
Table 2 demonstrates that CoCo-PINNs exhibit meaningfully better performance in the shapes of ‘fish’, ‘kite’, and ‘spike’ compared to classical PINNs. Although the credibility error for classical PINNs appears similar to that for CoCo-PINNs on the ‘square’ shape as shown in Table 2, the trained forward solution by classical PINNs illustrates an exorbitant large deviation that does not coincide with the analytic solution derived from the inverse parameter result, as shown in Figure 7. This indicates that, despite its strong performance in minimizing the loss function, the classical PINNs approach fails to effectively function as a forward solver.
Figure 7: Comparison of credibility: For the fish-shaped inclusion, the trained forward solution $u_{\text{NN}}$ obtained by classical PINNs and CoCo-PINNs is compared with the analytic solution $u_{p}$. The closer $u_{\text{NN}}$ approximates $u_{p}$, the more credible the method is considered to be.
4.4 Consistency of classical PINNs and CoCo-PINNs
In this subsection, we examine whether repeated experiments consistently yield similar results. Figure 8 shows the interface functions after training the classical PINNs and CoCo-PINNs, performed independently multiple times. We repetitively test $30$ times under the same condition and plot the interface function pointwise along the boundary of the unit disk. The blue-dashed and red-bold lines represent the mean of the interface functions produced by classical PINNs and CoCo-PINNs, respectively, while the blue- and red-shaded regions indicate the pointwise standard deviations of the interface functions, respectively. Each column corresponds to an experiment with ‘square’, ‘fish’, ‘kite’, and ‘spike’.
Figure 8: Consistency of interface functions: Repeated reconstructions of the interface function obtained by CoCo-PINNs and classical PINNs. The results demonstrate that CoCo-PINNs produce more consistent interface functions across trials compared to classical PINNs.
As shown in Figure 8, the interface functions trained with classical PINNs are consistent only for the spike-shaped example, whereas CoCo-PINNs achieve consistency across all examples. The precise value for the mean of the standard deviations is provided in column $3$ of Table 3. Additionally, columns $4$ through $6$ of Table 3 present the three quantities in 10, 11 and 12, which show the credibility and the errors indicating the neutral inclusion effect. These results clearly demonstrate that CoCo-PINNs exhibit superior credibility compared to classical PINNs. Furthermore, CoCo-PINNs outperform classical PINNs in achieving the neutral inclusion effect, particularly for complex shapes, as verified by the quantity $\|u_{p}-H\|_{\text{P-Neutral}}$. In some cases, the trained forward solution $u_{\text{NN}}$ and the inverse solution $p_{\text{NN}}$ produced by classical PINNs show a significant discrepancy, with $u_{\text{NN}}$ differing substantially from the analytic solution derived from $p_{\text{NN}}$ (see Figure 10 in Section 5.2).
In Figure 8, classical PINNs again yield nearly constant-valued solutions –just as they do for the disk– and therefore fail to effectively capture the problem for the shape with periodic spikes.
Table 3: Mean of the standard deviation of the interface function, and mean and standard deviations of errors for fitting the background fields after training. CoCo-PINNs show superior results to classical PINNs for complex shapes such as ‘fish’, ‘kite’, and ‘spike’. The ‘Mean of S.D.’ denotes the mean of the standard deviation for each point, and ‘S.D.’ denotes the standard deviation.
Shape Interface function $\|u^{\text{ext}}_{\text{NN}}-u_{p}\|_{\text{Cred}}$ $\|u^{\text{ext}}_{\text{NN}}-u_{p}\|_{\infty}$ $\|u_{p}-H_{u}\|_{\text{P-Neutral}}$
Mean of S.D. Mean S.D. Mean S.D. Mean S.D.
Classical PINNs square 2.828e-01 2.645e-03 4.383e-04 1.286e-01 2.013e-02 4.458e-03 6.529e-04
fish 9.217e-01 2.358e-03 2.787e-03 1.366e-01 3.835e-02 3.635e-03 4.211e-03
kite 1.199e+00 6.607e-02 3.327e-01 1.356e+00 3.573e+00 6.872e-02 3.485e-01
spike 9.038e-03 8.429e-03 8.331e-05 4.129e-01 3.989e-03 3.076e-04 2.372e-05
CoCo- PINNs square 1.013e-01 2.868e-03 9.634e-05 9.516e-02 2.278e-03 4.858e-03 1.682e-04
fish 6.646e-02 7.475e-04 3.037e-05 1.080e-01 2.383e-03 1.093e-03 4.563e-05
kite 1.009e-01 4.181e-04 8.149e-05 2.180e-01 2.505e-02 6.031e-04 1.356e-04
spike 7.068e-02 9.518e-03 3.681e-04 4.077e-01 6.375e-03 3.441e-04 1.440e-04
4.5 Stability of classical PINNs and CoCo-PINNs
In this subsection, we assess the stability of the interface function along with the change of environments of PDEs. Since both classical PINNs and CoCo-PINNs are trained for a fixed domain $\Omega$ and background field $H$, we focus on stability with respect to different conductivities $\sigma_{c}$. In Figure 9, we present the experimental results for various conductivity contrasts, where $\sigma_{c}=3,4,5,6,7$ and $\sigma_{m}=1$. The experiments were conducted for four different inclusion shapes: ‘square’, ‘fish’, ‘kite’, and ‘spike’. Recall that the ill-posed nature of inverse problems can lead to significant instability, causing the inverse solution to exhibit large deviations in response to environmental changes or re-experimentation. Classical PINNs for neutral inclusions with imperfect conditions represent such an unstable case, as demonstrated in Figure 9. In contrast, our CoCo-PINNs are stable for repeated experiments, and we confirmed that CoCo-PINNs are stable for slightly changed environments. Table 4 provides the mean of standard deviations for all experiments used in Figure 9. CoCo-PINNs demonstrate substantially improved consistency and stability over classical PINNs, as evidenced in Table 4.
Figure 9: Stability of interface functions under varying conductivities: The first and second rows display the interface functions obtained from classical PINNs and CoCo-PINNs, respectively. The mean values of the standard deviations are reported in Table 4. Across all cases, the interface functions learned by CoCo-PINNs exhibit significantly greater stability with respect to variations in conductivity compared to those from classical PINNs.
Table 4: Mean of standard deviations of interface functions for stability experiments.
Shape Conductivities for interior
$\sigma_{c}=3$ $\sigma_{c}=4$ $\sigma_{c}=5$ $\sigma_{c}=6$ $\sigma_{c}=7$
Classical PINNs square 2.679e-01 2.944e-01 2.828e-01 3.410e-01 3.316e-01
fish 9.699e-01 1.208e+00 9.217e-01 9.211e-01 7.013e-01
kite 1.278e+00 1.131e+00 1.199e+00 9.786e-01 9.539e-01
spike 1.281e-02 8.433e-03 9.038e-03 1.364e-02 7.030e-03
CoCo- PINNs square 1.200e-01 1.057e-01 1.013e-01 1.240e-01 1.176e-01
fish 6.505e-02 7.071e-02 6.646e-02 7.784e-02 7.521e-02
kite 1.139e-01 1.016e-01 1.009e-01 1.135e-01 1.153e-01
spike 2.379e-02 7.489e-02 7.068e-02 4.121e-02 5.800e-02
5 Further experiments
This section presents additional experimental results to assess the neutral inclusion effect, demonstrate the credibility for the four shapes defined in Figure 4, and compare the training time of both methods.
5.1 Neutral inclusion effect for arbitrary fields
@@
In this subsection, we present the neutral inclusion effect of shapes and different fields $H(x_{1},x_{2})=x_{1},x_{2}$, and $2x_{1}-x_{2}$. The neutral inclusion effects for one random experiment result are given in Table 5.
@@
CoCo- PINNs square 4.738e-03 4.981e-03 2.395e-02
fish 1.113e-03 2.568e-04 4.491e-03
kite 4.664e-04 3.839e-04 2.361e-03
spike 3.823e-04 3.858e-04 1.937e-03
After many fair experiments with both CoCo-PINNs and classical PINNs, we concluded that CoCo-PINNs were superior. After that, we tested the neutral inclusion effect for each shape by utilizing the CoCo-PINNs. Figure 6 presents the results. In the case of unit circle inclusion, exact neutral inclusion appeared. The shape we used may have no interface functions that make exact neutral inclusions. Notwithstanding, the CoCo-PINNs results for the neutral inclusion effect are, to some extent, satisfactory.
5.2 Illustrations for credibility
All experiments described in Table 2 are illustrated in Figures 11 and 10, by utilizing CoCo-PINNs and classical PINNs, respectively. We illustrate the pairs of
@@ 5.2 Illustrations for credibility
for each shape in Figure 4.
Figure 10: Experiment results using classical PINNs: For kite and spike shapes, notable discrepancies are observed between the forward solutions obtained via classical PINNs and the corresponding analytic solutions.
Figure 11: Experiment results using CoCo-PINNs: For all shapes considered, the forward solutions obtained via CoCo-PINNs more closely match the analytic solutions than those obtained by classical PINNs.
As shown Figures 11 and 10, the trained forward solutions by CoCo-PINNs $u_{\text{NN}}$ are more similar to the analytic solution $u_{p}$ using the interface parameter given by the CoCo-PINNs’ training results than the classical PINNs’ one.
5.3 Training time
Table 6: Training time.
Shape Time (s): Mean of 30 experiments
Classical PINNs square 1259.2729
fish 1254.7109
kite 1260.5721
spike 1257.5664
CoCo- PINNs square 1670.5210
fish 1671.7479
kite 1667.0469
spike 1668.4443
Table 6 presents a simple comparison of two methods in terms of computation time. We use a CPU (Intel® Xeon® Gold 6530) and a GPU (NVIDIA RTX 6000 Ada Generation). The learning time for CoCo-PINNs’ results is approximately 1.3 times that of classical PINNs’ results. The reported training times are averaged over 30 repeated experiments.
5.4 Ablation experiments on the regularization term for CoCo-PINNs
We examined the impact of the regularization term in the loss function in 9. The ablation experiments demonstrate that regularization generally enhances the consistency of the recovered interface function $p$ for CoCo-PINNs. Specifically, the regularized model exhibits clearly reduced variability compared to the unregularized model (see Figure 12). Also, in experiments for spike geometries, we observed that the regularized model maintains more consistent periodicity and amplitude in the oscillatory interface function (see also Figure 12).
Figure 12: Consistency comparison of CoCo-PINNs with and without regularization for fish- and spike-shaped inclusion.
6 Conclusion
We focus on the inverse problem of identifying an imperfect function that makes a given simply connected inclusion a neutral inclusion. We introduce a novel approach of Conformal mapping Coordinates Physics-Informed Neural Networks (CoCo-PINNs) based on complex analysis and PDEs. Our proposed approach of CoCo-PINNs solves both the forward and inverse problems simultaneously, much more effectively than the classical PINNs approach. While the classical PINNs approach may occasionally demonstrate success in finding an imperfect function with a strong neutral inclusion effect, the reliability of this performance remains uncertain. In contrast, CoCo-PINNs present high credibility, consistency, and stability, with the additional advantage of being explainable through analytical results. The potential applications of this method extend to analyzing and manipulating the interaction of embedded inhomogeneities and surrounding media, such as finding inclusions having uniform fields in their interiors. Several questions remain, including the generalization to multiple inclusions and three-dimensional problems, as well as proving the existence of an interface function that achieves neutrality.
The primary motivation for using conformal mapping is to establish a coordinate system that facilitates the automatic generation and classification of collocation points. Furthermore, employing these coordinates allows for a direct comparison between our numerical solution and the analytic solution, which is formulated in conformal mapping coordinates. It is worth noting that conformal maps may induce length distortion; for shapes causing severe distortion, an adaptive collocation strategy may be necessary to maintain accuracy. While conformal mappings generally do not exist for arbitrary shapes in three dimensions, the core strategy–establishing a method that naturally defines normal directions and distinct interior/exterior domains–remains applicable. One potential approach is the use of machine learning methods, such as Deep Signed Distance Function (DeepSDF) models [48].
Appendix A Geometric function theory
Geometric function theory is the research area of mathematics with the corresponding geometric properties of analytic functions. One remarkable result is the Riemann mapping theorem. We briefly introduce this theorem with related results. A connected open set in the complex plane is called a domain. We say that a domain $\Omega$ is simply connected if its complement $\mathbb{C}\setminus\Omega$ is connected.
Theorem A.1 (Riemann mapping theorem).
$\Psi(w)=w+a_{0}+\frac{a_{1}}{w}+\frac{a_{2}}{w^{2}}+\cdots.$ (17)
The quantity $\gamma$ in 17 is called the conformal radius of $\Omega$. One can obtain 17 by using Theorem A.1, the power series expansion of an analytic function, and its reflection with respect to a circle; we refer to for instance [47, Chapter 1.2] for the derivation.
We further assume, in accordance with the assumptions in [45], that $\Omega$ has an analytic boundary, that is, $\Psi$ can be conformally extended to $w\in\mathbb{C}:|w|>\gamma-\epsilon$} for some $\epsilon>0$.
Acknowledgement
This work is supported by the National Research Foundation of Korea(NRF) grants funded by the Korea government(MSIT) (RS-2024-00359109, RS-2025-02303239). This work is also supported by the Institute of Information & Communications Technology Planning & Evaluation (IITP) grant funded by the Korea government (MSIT) [RS-2021-II211341, Artificial Intelligence Graduate School Program (Chung-Ang University)].